part B1. The apparatus shown below consists of three 1.00L bulbs connected by stopcocks and held at the temperatures shown. Bulb A contains a mixture of H2O, CO2, and N2 at 25°C and a total pressure of 564 mmHg. Bulbs B and C are empty. CO2 sublimes at -78°C and N2 boils at -196°C 25 °C -70。 -190 °C a. The stopcock between A and B is opened. The pressure in A and B is now 219 mmlHg. What do bulbs A and B contain? (5 points) b. How many moles of H20 are in the system(5 points) c. Both stopcocks are opened. The pressure throughout the system is 33.5mmHg. What do bulbs A, B and C contain? (10 points) d. How many moles of N2 are in the system? (10 points) c. How many moles of CO2 are in the system? (10 points)
Answer #1

a. at - 70 c , H2o(g) converts into ice = H2o(s), other two are gases only.

no of mol of gas in A before openining stopcock

         = PV/RT

         = ((564/760)*1)/(0.0821*298)

         = 0.0303 mol

after openining stopcock

no of mol of gas mixture = PV/RT

         = ((219/760)*1)/(0.0821*298)

         = 0.012 mol

b) no of mol of water = 0.0303 - 0.012 = 0.0183 mol

c) after openining both stopcocks,

    bulb A contains = water,N2

    bulb B contains = water, N2

    bulb C, contains = CO2, N2

D) No of mol of N2 = PV/RT

         = ((33.5/760)*1)/(0.0821*298)

        = 0.0018 mol

c) No of mol of CO2 in the system = 0.0303- (0.012+0.0018) = 0.0165 mol